A spring with a spring constant of 1200 N/m has a 55-g ball at its end. The energy of the system is 5.0 J .

a) What is the amplitude A of vibration?

b) What is the maximum speed of the ball?

c) What is the speed when the ball is at a position x=+A/2.
Answer #1

Solution:

a) Using E = 0.5 k * A^2

=> 5 = 0.5 * 1200 * A^2

=> A = 9.1 cm

b) Using E = 0.5 mv^2

=> 5 = 0.5 * 0.055 v^2

=> v = 13.48 cm

c) Using conservation of energy

=> 5 = 0.5 * 1200 * (9.1/2)^2 + 0.5 * 0.055 * v^2

=> v = 11.7 m/s

thanks.